OpenStudy (anonymous):

1/64m^3+n^3 Factor the following expression please please help (it my last set of prob. on my hw)

OpenStudy (anonymous):

you have the sum of two cubes 1/4 m and n are the two cubes. Do you know the formula for sum of cubes?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[a^3+b^3=(a+b)(a^2-ab+b^2)\] and in this case a=1/4m and b=n

OpenStudy (anonymous):

1/4m or 1/64?? i am not seeing where u get 1/4??

OpenStudy (anonymous):

the cube root of 1/64 is 1/4 \[(1/4)^3 = 1/64\]

OpenStudy (anonymous):

oh ok sorry i get it

OpenStudy (anonymous):

if you type your answer, I will let you know if you are correct!

OpenStudy (anonymous):

1/4^3+n^2=(1/4^3+n)(1/4^2+1/4n+n^2)

OpenStudy (anonymous):

what happened to m :( (1/4 m+n)(1/16 m^2-1/4mn+n^2) a=1/4 m b=n a^2 = 1/16m^2 ab = 1/4 mn b^2 = n^2

OpenStudy (anonymous):

thanks. I realize my mistake when i see them so that is good. I am in a struggle with this math class this sem. I keep frogeting simple things.Thanks for your help.

OpenStudy (anonymous):

practice and you will get better!!!! Smart choice choosing to come here!